Minder Industries stock has a beta of 1.25. The company just paid a dividend of $.40, and the dividends are expected to grow at 5 percent. The expected return on the market is 12 percent, and Treasury bills are yielding 6.4 percent. The most recent stock price for the company is $81.

a.

Calculate the cost of equity using the DCF method. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. Calculate the cost of equity using the SML method. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

a.DCF method

b.SML method
Answer #1

a.

Using DCF Model,

Stock Price = D0(1 + g)/(r - g)

81 = 0.40(1.05)/(r - 0.05)

r = 0.40(1.05)/81 + 0.05

r = 0.0552

Cost of Equity = 5.52%

b.

Using SML Method,

Cost of Equity = Rf + Beta(Rm - Rf)

Cost of Equity = 0.064 + 1.25(0.12 - 0.064)

Cost of Equity = 0.134

Cost of Equity = 13.40%